OpenStudy (anonymous):

Find the distance between the points (-4,15) and (8,-1). If necessary, round you answer to tow decimal places. A) 10.58 units B) 14 units C) 20 units D) 5.29 units

OpenStudy (lgbasallote):

\[\Large \sqrt{[8 - (-4)]^2 + [-1 - 15]^2}\] \[\Large \sqrt{12^2 + (-16)^2}\] \[\Large \sqrt{144 + 256}\] \[\Large \sqrt{400}\] do you get it?

OpenStudy (anonymous):

so would it be 20?

OpenStudy (lgbasallote):

i used the distance formula \[\Large distance = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] and yes 20

OpenStudy (anonymous):

Thank you

OpenStudy (lgbasallote):

\[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]